OpenStudy (anonymous):

related rates problem need help stat! please!! Two cars start moving from the same poin. one travels south at 60mi/h and the other travels west at 25 mi/h. at what rate is the distance between the cars increasing two hours later?

OpenStudy (anonymous):

\[d(t)=\sqrt{3600t^t+625t^2}=\sqrt{4225t^2}\]Differentiate with respect to t, and you are done.

OpenStudy (anonymous):

By my best estimate, it's about 65 mph.

OpenStudy (anonymous):

Thank you so much !! could you also help me out on another related rates problem that I just posted I would greatly appreciate it !!

OpenStudy (anonymous):

No sweat. Do math every day.

OpenStudy (anonymous):

the problem reads; The top of a ladder slides down a verticle wall at a rate of 0.15m/s. at the moment when the bottom of the ladder is 3m from the wall, it slides away from the wall at a rate of 0.2 m/s. how long is the ladder?

OpenStudy (anonymous):

Thank you!